Operations Report – November 29th

SOL: 2

Name of person filing report: Ashley Kowalski

Non-nominal systems: (1) Anomaly with the front air lock, (2) “Push” button on EVA suit #7 is broken.

Notes on non-nominal systems: Upon returning from our EVAs, we noticed that the front air lock was difficult to close. The crew members needed to almost lift the door in order to be able to close it. It seemed as though the lower hinge was broken. The crew contacted Tier 1 support regarding this issue. The crew is also willing and interested in scheduling an EVA to try to fix this issue, if mission support is in agreement and the necessary tools/parts are available to us.

Regarding the broken “push” button in suit #7, the crew used a few tools from the RAM and made attempts to fix this, but were unsuccessful. Tier 1 support was contacted, and we are awaiting further instructions on how to fix this. Please note that the broken “push” button does not seem to affect the ability for the suit to charge.


Spirit rover used: Yes
Hours: 201.7 (Start), 201.8 (End)
Beginning charge: 100%
Ending charge: 88%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 105.6 (Start), 105.7 (End)
Beginning charge: 100%
Ending charge: 90%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 214.4 (Start), 214.5 (End)
Beginning charge: 100%
Ending charge: 91%
Currently charging: Yes

Perseverance rover used: Yes
Hours: 250.1 (Start), 250.3 (End)
Beginning charge: 100%
Ending charge: 85%
Currently charging: No

General notes on rovers: Three of the four rovers were left charging after the EVAs today. The crew engineer reminded all crew members that that is not the correct protocol. We also received this reminder message from Tier 1 support. The rovers will be left not charging after future EVAs.

Summary of Hab operations:
WATER USE: 20 gallons used today, 45 gallons used in total thus far
Water (static tank): 505 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios: The two walkie talkies mentioned in yesterday’s report (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: GreenHab Officer received training today. Additional information can be found in the GreenHab Report.
WATER USE: ~15 gallons
Heater: On
Supplemental light: Off
Harvest: No crops harvested.

Summary of ScienceDome operations: The crew set up their microscope in the ScienceDome today after all of the EVAs were completed. Two extension cords that the crew found in the RAM were used for this set up.

Dual split: No heat or AC used. All systems off.

Summary of RAM operations: Pliers were taken from the RAM into the habitat in order to attempt to fix the “push” button on EVA suit #7. Two extension cords were taken from the RAM in order to power the microscope and supplemental microscope light.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: None.

Copyright © The Mars Society. All rights reserved. | Main Site